Tylerr:

Solve this A salesman sold twice as much pears in the afternoon than in the morning. If he sold 360 kilograms of pears that day, how many kilograms did he sell in the morning and how many in the afternoon?

Hoodmemes:

For rn u wanna focus on this part "sold twice as much pears in the afternoon than in the morning." So he sold what ever kilo of pears in the morning but (2x) that many kilo's of pears in the afternoon. So I think the equation u would use to solve this would look like this: \[x + 2x = 360 \] (x) representing the kilo's of pears sold in the morning,(2x) representing the kilo's of pears sold in the morning but times (2),and 360 representing the total amount of kilo's of pears sold that day. \[x+2x=3x \] So u should (Divide) (360) kilo's of pears by (3x which also can just be written as 3) \[360\div3=120\] Now we have already found how many kilo's were sold in the morning,as stated in the problem "sold twice as much pears in the afternoon than in the morning" Basically u just wanna (Multiply) (120) which is the amount of kilo's sold in the morning by (2) \[120\times2=240\] (We can quickly check if its correct) \[120+240=360\] So (120) kilograms of pears were sold in the (Morning),and (240) kilograms of pears were sold in the (Afternoon) (Btw why isn't this posted in math?)

Tylerr:

Wow

Tylerr:

That’s a lot to read

Hoodmemes:

Just wanna know if u understand this?

Tylerr:

I don’t XD

Hoodmemes:

Great glad i could help ya

Tylerr:

lol thx!